OpenStudy (anonymous):

Use the commutative law of addition to write an equivalent expression 5a+b

OpenStudy (anonymous):

commutative : a+b=b+a so for this 5a+b=b+5a

OpenStudy (anonymous):

exact

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

anytym ! :-)